If the supply equation is given as Qs = -200+10P, the price elasticity of supply between R40 and R50, calculated using the arc elasticity of demand is
1
Expert's answer
2018-01-29T10:05:07-0500
If Qs = -200 + 10P, P1 = R40, P2 = R50, Q1 = -200 + 10*40 = 200, Q2 = -200 + 10*50 = 300, then the price elasticity of supply between R40 and R50 is Es = (300 - 200)/(50 - 40)*(50 + 40)/(300 + 200) = 100/10*90/500 = 1.8, so the supply is elastic.